# Constraining cosmological parameters using density split lensing and the conditional stellar mass function

### v1

@@ VII Estimating a covariance matrix
To measure a DESI-like SMF covariance from the T17, we first populate the haloes with the description in Sect. IV, and the reference parameters shown as the orange dot in Fig. 6 and given Table 1. Although we know that data release one of DESI will not cover the full UNIONS footprint, we assume it for this work and use the same masking we used for the UNIONS mocks, which we determined from the UNIONS galaxies themselves. Therefore, the parameter forecasts in this paper are over-optimistic for at least the first data releases of DESI. Given the stellar mass limit in Eq. V.2, we next measure the SMF for $1210.5<\log_{10}(M_{\star}/(h^{-1}\mathrm{M}_{\odot}))<$ 12 with $\Delta\log_{10}(M_{\star}/(h^{-1}\mathrm{M}_{\odot}))=0.1$ and $0.40.1<z<$ 0.4.
@@ VIII Building the emulator
To build the emulator, we closely follow the procedure of [4], with the exception that we are not averaging multiple noise seeds of the HOD, as the sample size for the used BGS mocks is much greater than the LOWZ sample used in [4], and the measurement of the ESD is more time-consuming. However, we do correct for cosmic variance by averaging over the $n_{\mathrm{p}}=25$ phases. For a given HOD h the corrected model is given by
[…]
@@ IX Forecast
where in the last step we assumed that $C^{\mathrm{emu}}\ll C^{\mathrm{T17}}$. This would, in principle, allow us to apply the correction in Eq. VIII.4 only to the $C^{\mathrm{T17}}$ and leaving $C^{\mathrm{emu}}$ untouched, which should improve the constraining power. However, $C^{\mathrm{emu}}\approx C^{\mathrm{T17}}$ for some scales, which did not allow us to apply that method. A completely alternative approach to incorporate the error of the emulator would be to use a Bayesian neural network, which finds a likelihood distribution for the parameters for each hidden node of a neural network instead of its best values. This allows us to get a sample of predictions for each input parameter, from which the mean and its uncertainty can be measured. We have tried this but did not achieve a similar accuracy as with CosmoPower, so we left it for future analysis to explore it in more detail.
@@
Pseudo simulation-based inference
We investigated a simulation-based inference (SBI) approach as an alternative to the MCMC approach. The general idea of SBI is to perform a parameter forecast if the likelihood is not well approximated with a Student’s $t$ or a Gaussian distribution [52, 53]. In other words, we do not approximate the likelihood as given in Eq. (IX.1) and then use an MCMC approach to sample this likelihood, but instead use a machine learning approach to learn the likelihood or posterior directly. In these works, the measured summary statistics $\mathbf{t}$ vary with parameters $\Theta$, where the $\Theta$ include cosmological, astrophysical and nuisance parameters. Moreover, one needs to incorporate the noise of the data into the measured $\mathbf{t}$ (i.e. shape and shape noise). The posterior is then estimated by minimizing the Kullback-Leibler divergence [54] between the parametric density estimator $p(\Theta|\mathbf{t};\mathbf{w})$, where $\mathbf{w}$ are the weights of the neural network, and the target distribution $p^{*}(\Theta|\mathbf{t})$:
$D_{\mathrm{KL}}(p^{*}|p)=\int p^{*}(\Theta|\mathbf{t})\ \ln\left(\frac{p(\Theta|\mathbf{t};\mathbf{w})}{p^{*}(\Theta|\mathbf{t})}\right)\ \mathrm{d}\mathbf{t}\ .$ (IX.5)
$-\ln U(\mathbf{w}|\mathbf{t},\Theta)=-\sum_{i=1}^{N_{\mathrm{samples}}}\ln[p(\Theta_{i}|\mathbf{t}_{i};\mathbf{w})]\,,$ (IX.6)
  1. 1.
    Use the emulator to predict a noise-free $\mathbf{t}$;
  2. 2.
    draw a multivariate random vector $G(0,C^{\mathrm{T17}})$, where $C^{\mathrm{T17}}$ is the T17 covariance matrix, and add it to $\mathbf{t}$;
  3. 3.
    draw a multivariate random vector $G(0,C^{\mathrm{emu}})$, where $C^{\mathrm{emu}}$ is the emulator covariance matrix, and add it to the already noisy $\mathbf{t}$.
We call this a pseudo-SBI, as we add the noise to the data using a multivariate Gaussian distribution instead of directly incorporating it into the simulations. Compared to the MCMC approach, this pseudo-SBI approach has the advantage of adding the noise coming from the data and the inaccuracies from the emulator separately to the measurements $\mathbf{t}$. If we want to be more similar to the MCMC approach, we should create the $\mathbf{t}$ measurements like
  1. 1.
    Use the emulator to predict a noise-free $\mathbf{t}$;
  2. 2.
To derive the posteriors we use the sbi [^10] https://github.com/sbi-dev/sbi package [55], which is the successor of pydelfi [56]. In particular, we are using the neural posterior estimator version A (NPE-A), which uses Bayesian Conditional Density Estimation [57] to estimate the posterior, which in turn uses a mixture density network [58]. We provided a flat prior during the training and removed all points of the chain after evaluating $p(\Theta|\mathbf{t}=\mathbf{t}_{0};\mathbf{w})$, that fell outside our non-flat prior given in Fig. 6. The reference $\mathbf{t}_{0}$ is the noise-free prediction of the emulator at the parameters given in Table 1.
In Fig. 13, we show the two approaches of incorporating noise to the $\mathbf{t}$ are almost identical. Comparing the MCMC approach to the SBI, we see that the pseudo-SBI posterior is slightly enlarged. Although the pseudo-SBI gives slightly larger constraints, we want to emphasize the match, especially in the cosmological parameters. Furthermore, it is worth mentioning that the pseudo-SBI produces converged posteriors roughly five times faster than emcee, depending on the number of walkers and steps of the MCMC process and the measurements to train to the neural network of the SBI.
Figure 13: Posterior distribution using the pseudo-SBI approach, where we use two different approaches to incorporate the noise in into the measured summary statistics $\mathbf{t}$.
@@ Acknowledgements.
We thank Enrique Paillas, Hanyu Zhang, and Marco Bonici for fruitful comments regarding AbacusSummit and other analysis details. We want to thank Mike Jarvis for developing and maintaining treecorr, and Alessio Mancini for developing CosmoPower emulator, which provided a significant part of our model. Furthermore, we thank the authors of AbacusSummit and AbacusHOD for providing and maintaining these excellent simulations and code. We thank the UNIONS team, for providing us with preliminary shear catalogues and redshift distributions, which allowed us to perform a realistic parameter forecast. We acknowledge financial support from the Canadian Space Agency (Grant 23EXPROSS1), the Waterloo Centre for Astrophysics and the National Science and Engineering Research Council Discovery Grants program. This research was enabled in part by support provided by Calcul Quebec [^11] https://docs.alliancecan.ca/wiki/Narval and Compute Ontario [^12] https://docs.alliancecan.ca/wiki/Graham and the Digital Research Alliance of Canada131313alliancecan.ca. This research used resources from the National Energy Research Scientific Computing Center, supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231.
Appendix A T17 analysis
@@
In order to still perform a T17 analysis, we use the mean over all 756 measurements as our reference data vector for the forecast analysis. We would not necessarily expect an excellent fit due to the differences in mass resolutions and halo finder. Furthermore, T17 assumes that the neutrinos are massless, but the AbacusSummit have used $\omega_{\mathrm{ncdm}}=0.00064420$. This can have an $\sim 3\%$ effect on the matter power spectrum, especially on smaller scales [61], shifting the $\sigma_{8}$ up by $1.8\%$. However, since the AbacusSummit is not necessarily a better fit to real data, using the T17 simulations as a reference might indicate possible biases in real data analyses. In Fig. 14 and 15, we show the pyccl and emulator predictions compared to T17 measurements that are used as the reference data vector. We can see that, especially for higher stellar mass bins, the match of the ESD at small scales is less accurate.
If we consider only scales above $>8\,h^{-1}\mathrm{Mpc}$ the bias in the cosmological parameters is gone. The remaining inaccuracies at large scales are then compensated by adjusting especially the HOD parameters $B_{\mathrm{cen}}$ and $\sigma_{\mathrm{c}}$. Overall, we are not worried about modelling our summary statistics since these biases are at a 2-$\sigma$ level in high dimensional parameter space, and it is not clear which of the two simulations is the better match to real data.

### v2

@@ VII Estimating a covariance matrix
To measure a DESI-like SMF covariance from the T17, we first populate the haloes with the description in Sect. IV, and the reference parameters shown as the orange dot in Fig. 6 and given Table 1. Although we know that data release one of DESI will not cover the full UNIONS footprint, we assume it for this work and use the same masking we used for the UNIONS mocks, which we determined from the UNIONS galaxies themselves. Therefore, the parameter forecasts in this paper are over-optimistic for at least the first data releases of DESI. Given the stellar mass limit in Eq. V.2, we next measure the SMF for $10.5<\log_{10}(M_{\star}/(h^{-1}\mathrm{M}_{\odot}))<12$ with $\Delta\log_{10}(M_{\star}/(h^{-1}\mathrm{M}_{\odot}))=0.1$ and $0.1<z<0.4$.
@@ VIII Building the emulator
To build the emulator, we closely follow the procedure of [4], with the exception that we are not averaging multiple noise seeds of the HOD, as the sample size for the used BGS mocks is much greater than the LOWZ sample used in [4], and the measurement of the ESD is more time-consuming. However, to reduce the effect of cosmic variance, we average over the $n_{\mathrm{p}}=25$ phases. For a given HOD h, the corrected model is given by
$\bm{m}^{\mathrm{cor}}(\Theta;\texttt{h})=\frac{\bm{m}^{\texttt{ph00}}(\Theta)}{\bm{m}^{\texttt{ph00}}(\texttt{c000};\texttt{h})}\frac{1}{n_{\mathrm{p}}}\sum_{p=1}^{n_{\mathrm{p}}}\bm{m}^{p}(\texttt{c000};\texttt{h})\,\,$, (VIII.1)
@@ IX Forecast
where in the last step we assumed that $C^{\mathrm{emu}}\ll C^{\mathrm{T17}}$. This would, in principle, allow us to apply the correction in Eq. VIII.4 only to the $C^{\mathrm{T17}}$ and leaving $C^{\mathrm{emu}}$ untouched, which should improve the constraining power. However, $C^{\mathrm{emu}}\approx C^{\mathrm{T17}}$ for some scales, which did not allow us to apply that method. A completely alternative approach to incorporate the error of the emulator would be to use a Bayesian neural network, which finds a likelihood distribution for the parameters for each hidden node of a neural network instead of its best values. This allows us to get a sample of predictions for each input parameter, from which the mean and its uncertainty can be measured. We have tried this but did not achieve a similar accuracy as with CosmoPower [^9] https://alessiospuriomancini.github.io/cosmopower/, so we left it for future analysis to explore it in more detail.
@@ Acknowledgements.
We thank Enrique Paillas, Hanyu Zhang, and Marco Bonici for fruitful comments regarding AbacusSummit and other analysis details. We want to thank Mike Jarvis for developing and maintaining treecorr, and Alessio Spurio Mancini for developing CosmoPower emulator, which provided a significant part of our model. Furthermore, we thank the authors of AbacusSummit and AbacusHOD for providing and maintaining these excellent simulations and code. We thank the UNIONS team, for providing us with preliminary shear catalogues and redshift distributions, which allowed us to perform a realistic parameter forecast. We acknowledge financial support from the Canadian Space Agency (Grant 23EXPROSS1), the Waterloo Centre for Astrophysics and the National Science and Engineering Research Council Discovery Grants program. This research was enabled in part by support provided by Calcul Quebec [^11] https://docs.alliancecan.ca/wiki/Narval and Compute Ontario [^12] https://docs.alliancecan.ca/wiki/Graham and the Digital Research Alliance of Canada131313alliancecan.ca. This research used resources from the National Energy Research Scientific Computing Center, supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231.
Appendix A Pseudo simulation-based inference
In this section, we investigated an alternative sampling strategy to the standard MCMC approach, which is based on the method of simulation-based inference (SBI). The general idea of SBI is to perform a parameter forecast if the likelihood is not well approximated with a Student’s $t$ or a Gaussian distribution [55, 56]. In other words, we do not approximate the likelihood as given in Eq. (IX.1) and then use an MCMC approach to sample this likelihood, but instead use a machine learning approach to learn the likelihood or posterior directly. In these works, the measured summary statistics $\mathbf{t}$ vary with parameters $\Theta$, where the $\Theta$ include cosmological, astrophysical and nuisance parameters. Moreover, one needs to incorporate the noise of the data into the measured $\mathbf{t}$ (i.e. shape and shape noise). The posterior is then estimated by minimizing the Kullback-Leibler divergence [57] between the parametric density estimator $p(\Theta|\mathbf{t};\mathbf{w})$, where $\mathbf{w}$ are the weights of the neural network, and the target distribution $p^{*}(\Theta|\mathbf{t})$:
$D_{\mathrm{KL}}(p^{*}|p)=\int p^{*}(\Theta|\mathbf{t})\ \ln\left(\frac{p(\Theta|\mathbf{t};\mathbf{w})}{p^{*}(\Theta|\mathbf{t})}\right)\ \mathrm{d}\mathbf{t}\ .$ (A.1)
$\ln U(\mathbf{w}|\mathbf{t},\Theta)=-\sum_{i=1}^{N_{\mathrm{samples}}}\ln[p(\Theta_{i}|\mathbf{t}_{i};\mathbf{w})]\,$, (A.2)
  1. 1.
    Use the emulator to predict a noise-free $\mathbf{t}$;
  2. 2.
We call this a pseudo-SBI, as we add the noise to the data using a multivariate Gaussian distribution instead of directly incorporating it into the simulations.
To derive the posteriors we use the sbi [^14] https://github.com/sbi-dev/sbi package [58], which is the successor of pydelfi [59]. In particular, we are using the neural posterior estimator version A (NPE-A), which uses Bayesian Conditional Density Estimation [60] to estimate the posterior, which in turn uses a mixture density network [61]. We provided a flat prior during the training and removed all points of the chain after evaluating $p(\Theta|\mathbf{t};\mathbf{w})$ at $\mathbf{t}=\mathbf{t}_{0}$, that fell outside our non-flat prior given in Fig. 6. The reference $\mathbf{t}_{0}$ is the noise-free prediction of the emulator at the parameters given in Table 1.
In Fig. 13, we compare the MCMC approach to the pseudo-SBI. Although we added Gaussian noise to the noise-free $\mathbf{t}$, we do not expect that both approaches result in identical posteriors. One reason is that standard MCMC and SBI are approximating the posterior differently. Another reason is that SBI seems to have smoother posteriors, which could mean that the SBI is smoothing over the remaining noise in the emulator. Therefore, we are unsurprised that the pseudo-SBI posterior results in a slightly enlarged posterior distribution. Although the pseudo-SBI gives slightly larger constraints, we want to emphasize the match, especially in the cosmological parameters. Furthermore, it is worth mentioning that the pseudo-SBI produces converged posteriors roughly five times faster than emcee, depending on the number of walkers and steps of the MCMC process and the measurements to train to the neural network of the SBI.
Figure 13: Posterior distribution using the pseudo-SBI approach and its comparison to the standard MCMC approach.
Appendix B T17 analysis
@@
In order to still perform a T17 analysis, we use the mean over all 756 measurements as our reference data vector for the forecast analysis shown in Fig. 16 for the different summary statistics. We would not necessarily expect an excellent fit due to the differences in mass resolutions and halo finder. Furthermore, T17 assumes that the neutrinos are massless, but the AbacusSummit have used $\omega_{\mathrm{ncdm}}=0.00064420$. This can have an $\sim 3\%$ effect on the matter power spectrum, especially on smaller scales [62], shifting the $\sigma_{8}$ up by $1.8\%$. However, since the AbacusSummit is not necessarily a better fit to real data, using the T17 simulations as a reference might indicate possible biases in real data analyses. In Fig. 14 and 15, we show the pyccl and emulator predictions compared to T17 measurements that are used as the reference data vector. We can see that, especially for higher stellar mass bins, the match of the ESD at small scales is less accurate.
If we consider only scales above $8\,h^{-1}\mathrm{Mpc}$ the bias in the cosmological parameters is gone. The remaining inaccuracies at large scales are then compensated by adjusting especially the HOD parameters $B_{\mathrm{cen}}$ and $\sigma_{\mathrm{c}}$. Overall, we are not worried about modelling our summary statistics since these biases are at a 2-$\sigma$ level in high dimensional parameter space, and it is not clear which of the two simulations is the better match to real data.